OpenStudy (anonymous):

Please help if you can! solve the system by substitution. 2x+y=-11 3x-4y=11 a.(3,5) b.(-5,-3) c. (-3,-5) d.(5,3)

OpenStudy (tfht):

well take the first equation, and get y alone. so to do that, you subtract 2x from both sides to get y=-2x-11

OpenStudy (tfht):

then sub that in in the second equation for y to get (3x)-4(-2x-11)=11

OpenStudy (tfht):

then you get 3x+8x+44=11

OpenStudy (anonymous):

c. (-3,-5

OpenStudy (tfht):

then simplify to 11x=-33 so x=-3

OpenStudy (anonymous):

Thank you @tfht for explaining it.

OpenStudy (yttrium):

you have 2x+y = -11 compute for x or y using second equation let say solving for x 3x-4y = 11 3x = 4y+11 x = (4y+11)/3 then substitute x in the 1st equation. 2((4y+11)/3)+y = 11 then solve for y.

OpenStudy (tfht):

then you sub -3 in for x to get 3(-3)-4y=11 then simplify it to -9-4y=11 then -4y=20 so y=-5. and yup no problem! i love math and so i come on here when im bored lol. glad i could help!

OpenStudy (anonymous):

I don't understand it. but when people explain things to me it actually helps me understand it more. Math is not my strongest subject.

OpenStudy (anonymous):

@tfht can you help me with another problem?

OpenStudy (tfht):

sure!

OpenStudy (anonymous):

what is the solution of the following system. -3x-2y=-12 9x+6y=-9 (2,1) b.no solutions c. (-2,-1) d. infinitely many solutions

OpenStudy (anonymous):

no solution

OpenStudy (tfht):

ok so here i would simplify the second equation down to 3x+2y=-3

OpenStudy (tfht):

so then you could get x alone or y alone, im going to go for x and get 3x=-2y-3 then divide everything by 3 to get x=-2/3y-1 and sub that in for x in the first equation

OpenStudy (anonymous):

you will see nooooo solution

OpenStudy (tfht):

-3(-2/3y-1)-2y=-12 which equals 2y-2y=-12 so that simplifies down to 0=-12 but that isnt possible, so it's no solution

OpenStudy (anonymous):

Thank you to both of you! @tfht and @hazem-wardat

OpenStudy (tfht):

yup!

OpenStudy (anonymous):

wlc